OpenStudy (anonymous):

Simplify by removing factors of 1. 7q^2+7q/56q^2+49q

OpenStudy (anonymous):

q+1/8q+1 not sure that is right .

OpenStudy (anonymous):

=7*q(q+1)/7q(8q+7) 7*q simplyfy, so =(q+1)/(8q+7)

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

with the () correct

OpenStudy (anonymous):

what you mean?

OpenStudy (anonymous):

(q+1) like that

OpenStudy (anonymous):

or q+1

OpenStudy (anonymous):

if you dont put the braquest, the mathematical expression means different things. Put the braquets. makes everythin clearer

OpenStudy (anonymous):

understood thank you

OpenStudy (anonymous):

ok, uw

OpenStudy (anonymous):

Hey dude, Im not receiving any medal 4 the help! Dont foget it (that is the way of saying thanks in here)